One number exceeds another number by 26. the sum of these numbers is 64. What are these numbers? I need to know how to solve word problems like these, can you please walk me through the steps.
Answer by scott8148(6628)

"One number exceeds another number by 26" ___ n = m + 26

"the sum of these numbers is 64" ___ m + n = 64

substituting the "equivalent value" for n ___ m + (m + 26) = 64 ___ 2m + 26 = 64 ___ m = 19

substituting the "equivalent value" for m ___ n = (19) + 26 ___ n = 45